# Weighted Pseudorandom Generators for Read-Once Branching Programs via Weighted Pseudorandom Reductions

### v3

@@ 1.5 Technical Overview
Notice that this also immediately gives a WPRG for $f$ with error $\varepsilon_{0}^{k}(n+1)$ by definition. However the $k$ independent PRG callings cost too much randomness. [CDR+21] further reduces the randomness by using an INW generator to generate the seeds for the $k$ independent PRG callings. But since INW does not have optimal seed length, this takes $O(\log k\log\frac{kw}{\varepsilon}+s)$ randomness which has an extra $O(\log k)$ factor in the main term. Using the sampler technique of [Hoz21], one can avoid this factor, but only when the starting error $\varepsilon_{0}$ is already $1/\operatorname{\mathrm{poly}}w$. If the large error PRG has $\varepsilon_{0}\gg 1/\operatorname{\mathrm{poly}}w$, then this does not work as desired. For example, if the large-error PRG is the NZ PRG, then it is unclear how to use this to derive an optimal seed length for $\varepsilon=1/\operatorname{\mathrm{poly}}w$.
@@
More concretely, the $l$ length reductions mentioned above are instantiated by the Richardson iteration based error reduction with Armoni’s generator [Arm98, KNW08] as base PRGs. For a length-$n$, width-$w$ ROBP over $\Sigma=\{0,1\}^{s}$, we first choose $\varepsilon_{0}=2^{-\sqrt{\log w\log n}}$ and apply a $(d_{0}=\log K_{0},K_{0},\varepsilon)$-reduction from $\mathcal{B}_{n,s,w}$ to $\mathcal{B}_{n_{1},s^{\prime}_{1},w}$, in which $n_{1}=O\left(\sqrt{\frac{\log w}{\log n}}\right)$, $K_{0}=2^{O\left(\sqrt{\log n\log w}\right)}$ and $s^{\prime}_{1}=O\left(s+\frac{\log n\log(nw/\varepsilon_{0})}{\log\log w-\log\log(n/\varepsilon_{0})}\right)$. For the subsequent $l-1$ reductions i.e. for $i=1,2,\ldots,l$, we setup a $(d_{i}=\log K_{i},K_{i},\varepsilon_{i})$ reduction from $\mathcal{B}_{n_{i},s_{i},w}$ to $\mathcal{B}_{n_{i+1},s^{\prime}_{i+1},w}$, where we assume $s_{i}=c\log w$ with $c$ being a universal constant. We choose $\varepsilon_{i}=2^{-\log w/n_{i}^{1/3}}$, and consequently $n_{i+1}=n_{i}^{1/3}$, $K_{i}=O\left(2^{n_{i}^{1/3}\log n_{i}}\right)=o\left(\log^{1/2}w\right)$ and $s^{\prime}_{i+1}=O(s_{i}+\log w)$. We iterate for $l$ steps until we reach $n_{l}=O(1)$. One can see that $l\leq O(\log\log\log w)$.
@@
We start by reviewing the recent WPRG given by [CHL+23] in the view of weighted pseudorandom reductions. Their WPRG combines the INW generator and a new matrix iteration method. To fool a unbounded-width permutation ROBP of length $n$ and alphabet $\{0,1\}^{s}$ within error $\varepsilon$, they first prepare a base INW generator $\mathsf{PRG}:\{0,1\}^{d}\to\left(\{0,1\}^{s}\right)^{n}$ with moderate sv-error [^5] See Definition 4.5. $\tau=2^{O(\log\log n\sqrt{\log(1/\varepsilon)})}$. Then they construct the $\varepsilon$-WPRG $(G,\sigma)$ in the following form:
@@
To address this problem, we introduce the following new strategy based on weighted pseudorandom reductions. Let $\mathcal{P}_{n,s}$ denote the class of length-$n$ unbounded-width permutation ROBPs over alphabet $\{0,1\}^{s}$. Then $(G,\sigma)$ can be viewed as a $(\log K,K,\varepsilon)$-reduction from $\mathcal{P}_{n,s}$ to $\mathcal{P}_{m,d}$. Let our target error be $\varepsilon$. We setup $l$ reductions as the following. The $i$-th reduction is a $(d_{i}=\log K_{i},K_{i},\varepsilon_{i})$-reduction from $\mathcal{P}_{n_{i-1},s_{i-1}}$ to $\mathcal{P}_{n_{i},s_{i}}$. For the first reduction, we set $\tau_{1}=O\left(\sqrt{\log(1/\varepsilon)}\right)$ to obtain a $(\log K_{1},K_{1},\varepsilon/2)$-reduction from $\mathcal{P}_{n,s}$ to $\mathcal{P}_{n_{1},s_{1}}$. One can see $n_{1}=O\left(\log n\cdot\sqrt{\log(1/\varepsilon)}\right)$, $s_{1}=s+O\left(\log n\cdot\left(\log\log n+\sqrt{\log(1/\varepsilon)}\right)\right)$ and $K_{1}=\exp\left(O\left(\log n\cdot\sqrt{\log(1/\varepsilon)}\right)\right)$. For the subsequent $l-1$ reductions, we set $\varepsilon^{\prime}=(\varepsilon/(2K_{1}))^{2}$ and $\tau_{i}=2^{\frac{C\log(1/\varepsilon^{\prime})}{\log^{2}n_{i-1}}}$ to obtain a$(\log K_{i},K_{i},\varepsilon^{\prime})$-reduction from $\mathcal{P}_{n_{i-1},s_{i-1}}$ to $\mathcal{P}_{n_{i},s_{i}}$, where $C>0$ is a universal constant. One can deduce that $n_{i}=O\left(\log n_{i-1}\frac{\log(1/\varepsilon^{\prime})}{\log(1/\tau_{i})}\right)=\log^{3}n_{i-1}$, $s_{i}=s_{i-1}+O(\log n_{i-1}\cdot(\log\log n_{i-1}+\log(1/\tau_{i})))=s_{i-1}+O\left(\frac{\log(1/\varepsilon^{\prime})}{\log n_{i-1}}\right)$ and $K_{i}=2^{O(\log^{3}n_{i})}$. We do the iteration until $n_{l}=O(1)$. So $l=o(\log\log n)$.

### v4

@@ 1.5 Technical Overview
Notice that this also immediately gives a WPRG for $f$ with error $\varepsilon_{0}^{k/2}(n+1)$ by definition. However the $k$ independent PRG callings cost too much randomness. [CDR+21] further reduces the randomness by using an INW generator to generate the seeds for the $k$ independent PRG callings. But since INW does not have optimal seed length, this takes $O(\log k\log\frac{kw}{\varepsilon}+s)$ randomness which has an extra $O(\log k)$ factor in the main term. Using the sampler technique of [Hoz21], one can avoid this factor, but only when the starting error $\varepsilon_{0}$ is already $1/\operatorname{\mathrm{poly}}w$. If the large error PRG has $\varepsilon_{0}\gg 1/\operatorname{\mathrm{poly}}w$, then this does not work as desired. For example, if the large-error PRG is the NZ PRG, then it is unclear how to use this to derive an optimal seed length for $\varepsilon=1/\operatorname{\mathrm{poly}}w$.
@@
More concretely, the $l$ length reductions mentioned above are instantiated by the Richardson iteration based error reduction with Armoni’s generator [Arm98, KNW08] as base PRGs. For a length-$n$, width-$w$ ROBP over $\Sigma=\{0,1\}^{s}$, we first choose $\varepsilon_{0}=2^{-\sqrt{\log w\log n}}$ and apply a $(d_{0}=\log K_{0},K_{0},\varepsilon)$-reduction from $\mathcal{B}_{n,s,w}$ to $\mathcal{B}_{n_{1},s^{\prime}_{1},w}$, in which $n_{1}=O\left(\sqrt{\frac{\log w}{\log n}}\right)$, $K_{0}=2^{O\left(\sqrt{\log n\log w}\right)}$ and $s^{\prime}_{1}=O\left(s+\frac{\log n\log(nw/\varepsilon_{0})}{\log\log w-\log\log(n/\varepsilon_{0})}\right)$. For the subsequent $l-1$ reductions i.e. for $i=1,2,\ldots,l$, we setup a $(d_{i}=\log K_{i},K_{i},\varepsilon_{i})$ reduction from $\mathcal{B}_{n_{i},s_{i},w}$ to $\mathcal{B}_{n_{i+1},s^{\prime}_{i+1},w}$, where we assume $s_{i}=c\log w$ with $c$ being a universal constant. We choose $\varepsilon_{i}=2^{-(\log w)/n_{i}^{1/3}}$, and consequently $n_{i+1}=n_{i}^{1/3}$, $K_{i}=O\left(2^{n_{i}^{1/3}\log n_{i}}\right)=o\left(\log^{1/2}w\right)$ and $s^{\prime}_{i+1}=O(s_{i}+\log w)$. We iterate for $l$ steps until we reach $n_{l}=O(1)$. One can see that $l\leq O(\log\log\log w)$.
@@
We start by reviewing the recent WPRG given by [CHL+23] in the view of weighted pseudorandom reductions. Their WPRG combines the INW generator and a new matrix iteration method. To fool a unbounded-width permutation ROBP of length $n$ and alphabet $\{0,1\}^{s}$ within error $\varepsilon$, they first prepare a base INW generator $\mathsf{PRG}:\{0,1\}^{d}\to\left(\{0,1\}^{s}\right)^{n}$ with moderate sv-error [^5] See Definition 4.5. $\tau=2^{-O(\log\log n\sqrt{\log(1/\varepsilon)})}$. Then they construct the $\varepsilon$-WPRG $(G,\sigma)$ in the following form:
@@
To address this problem, we introduce the following new strategy based on weighted pseudorandom reductions. Let $\mathcal{P}_{n,s}$ denote the class of length-$n$ unbounded-width permutation ROBPs over alphabet $\{0,1\}^{s}$. Then $(G,\sigma)$ can be viewed as a $(\log K,K,\varepsilon)$-reduction from $\mathcal{P}_{n,s}$ to $\mathcal{P}_{m,d}$. Let our target error be $\varepsilon$. We setup $l$ reductions as the following. The $i$-th reduction is a $(d_{i}=\log K_{i},K_{i},\varepsilon_{i})$-reduction from $\mathcal{P}_{n_{i-1},s_{i-1}}$ to $\mathcal{P}_{n_{i},s_{i}}$. For the first reduction, we set $\tau_{1}=2^{-O\left(\sqrt{\log(1/\varepsilon)}\right)}$ to obtain a $(\log K_{1},K_{1},\varepsilon/2)$-reduction from $\mathcal{P}_{n,s}$ to $\mathcal{P}_{n_{1},s_{1}}$. One can see $n_{1}=O\left(\log n\cdot\sqrt{\log(1/\varepsilon)}\right)$, $s_{1}=s+O\left(\log n\cdot\left(\log\log n+\sqrt{\log(1/\varepsilon)}\right)\right)$ and $K_{1}=\exp\left(O\left(\log n\cdot\sqrt{\log(1/\varepsilon)}\right)\right)$. For the subsequent $l-1$ reductions, we set $\varepsilon^{\prime}=(\varepsilon/(2K_{1}))^{2}$ and $\tau_{i}=2^{-\frac{C\log(1/\varepsilon^{\prime})}{\log^{2}n_{i-1}}}$ to obtain a $(\log K_{i},K_{i},\varepsilon^{\prime})$-reduction from $\mathcal{P}_{n_{i-1},s_{i-1}}$ to $\mathcal{P}_{n_{i},s_{i}}$, where $C>0$ is a universal constant. One can deduce that $n_{i}=O\left(\log n_{i-1}\frac{\log(1/\varepsilon^{\prime})}{\log(1/\tau_{i})}\right)=\log^{3}n_{i-1}$, $s_{i}=s_{i-1}+O(\log n_{i-1}\cdot(\log\log n_{i-1}+\log(1/\tau_{i})))=s_{i-1}+O\left(\frac{\log(1/\varepsilon^{\prime})}{\log n_{i-1}}\right)$ and $K_{i}=2^{O(\log^{3}n_{i})}$. We do the iteration until $n_{l}=O(1)$. So $l=o(\log\log n)$.
@@ 5.2 The derandomization for regular ROBPs with large alphabets
Let $G=(U,V,E)$ be a $d$-regular bigraph. A consistent one-way labeling of $G$ is a one-way labeling of $G$ such that for every $v\in V$, the labels of the incoming edges of $v$ are distinct, i.e., the $G[v,i]=G[u,i]$ implies $u=v$. In this case, we can extend the labeling to a two-way labelings:
$\mathsf{Rot}_{G}(u,i)=(G[u,i],i),$
i.e., the incoming label and the outgoing label of an edge $(u,v)$ are the same.
Lemma D.3 (Derandomized Product with consistent consistent one-way labelings,[RV05]).
[…]
Let $f$ be a permutation ROBP in $\mathcal{P}_{(n,s)}$. Then $f$ is also a regular ROBP with a consistent labeling. It can be sv-approximated with the method of Lemma 5.12. Furthermore, the method corresponds to fooling $f$ with the INW generator.
Lemma D.4 (equivariance between INW generator and Lemma 5.12).
Let $f$ be a permutation ROBP in $\mathsf{P}_{(n,s)}$. Let $t\in[\log n]$, $(j,j+2^{t})\in E_{SC_{n}}$ and $\widetilde{G}_{j,j+2^{t}}$ be the $(d\cdot c^{t})$-regular bigraph with two-way labeling as defined in Lemma 5.12. Then for any $u\in V^{j}$ and $x\in\{0,1\}^{d+t\log c}$, we have
$\exists x^{\prime},\widetilde{G}_{j,j+2^{t}}[u,x]=[v,x^{\prime}]\Leftrightarrow\left[f^{[j,j+2^{t}]}(\mathsf{INW}(x))\right]_{u,x}=v,$
Proof.
The proof is by induction on $t$. For the base case $t=0$, both sides mean $u$ has an outgoing edge labeled $x$ going to $v$. Assume the statement holds for $t-1$. Recall that
[…]
Then for $t$, the left side can be computed as follows (where $x=(z,y)$):
$(w,z)=\mathsf{Rot}_{\widetilde{G}_{j,j+2^{t-1}}}(u,z),$
$(z^{\prime},y^{\prime})=\mathsf{Rot}_{H_{t}}(z,y),$
$(v,w)=\mathsf{Rot}_{\widetilde{G}_{j+2^{t-1},j+2^{t}}}(z^{\prime},y^{\prime}).$
Where the right side can be computed as follows:
$f^{[j,j+2^{t-1}]}(\mathsf{INW}_{t-1}(z))=w,$
$H_{t}[z,y]=z^{\prime},$
$f^{[j+2^{t-1},j+2^{t}]}(\mathsf{INW}_{t-1}(z^{\prime}))=v.$
The induction hypothesis implies that the two sides are equivalent. ∎
Finally, we can prove the lemma.
Lemma D.5 (Lemma 4.6 restated).
$s+O(\log n(\log\log n+\log(1/\varepsilon))).$
Proof.
Let $\lambda=\min\{\varepsilon/(11\log n),1/6(\log^{2}n)\}$. Let $d=2^{s}$ and for each $t\in[\log n]$, let $H_{t}$ be a $c$-regular expanders with $\lambda(H_{t})\leq\lambda$ from Lemma 5.14. Let $\{\mathsf{INW}_{t}\}_{t\in[\log n]}$ be the INW generator from the definition. Let $\mathsf{INW}=\mathsf{INW}_{\log n}$. Then $\mathsf{INW}$ is a PRG with seed lengths
$\log d+t\cdot\log c=s+O(\log n(\log\log n+\log(1/\varepsilon))).$
For any $f\in\mathcal{P}_{(n,s)}$, the derandomized product of $f$ coincides with the the output of $f\circ INW$, i.e. $\widetilde{\mathbf{W}}_{n\leftarrow 0}=E_{X}f^{[0,n]}(\mathsf{INW}(X))$. Since $\widetilde{\mathbf{W}}_{n\leftarrow 0}\overset{sv}{\approx}_{11\lambda\log n}\widetilde{\mathbf{W}}_{n\leftarrow n-1}\cdots\widetilde{\mathbf{W}}_{1\leftarrow 0}$ by Lemma 5.12, $INW$ fools $f$ with $\varepsilon$-sv-error. ∎
Appendix E Reduction for Regular ROBPs over Large Alphabets
In this section, we show that the WPRG construction in Section 3 of [CL24] naturally extends to the case of regular ROBPs with large alphabets. This will immediately give us Theorem 3.21. We start by introducing an equivalent definition of the weight of ROBPs,
Definition E.1 (An Equivalent Definition of Weight of ROBPs).
Let $f\in\mathcal{B}_{n,s,w}$ be an ROBP. Let $V=\bigcup_{i=0}^{n}V_{i}$ be its vertices and $E=\bigcup_{i=1}^{n}E_{i-1,i}$ be its edges, where $E_{i-1,i}\subseteq V_{i-1}\times V_{i}$ For every $0\leq l\leq r\leq n$, Let $\mathbf{M}_{l\ldots r}$ be the stochastic matrix of $f$ from layer $l$ to layer $r$. Let $e_{i}$ denote the unitary vector at the $i$-th coordinate. For every $y\in\mathbb{R}^{w}$, $i\in[n]$, define the layer $i$ weight of $f$ on $y$ as:
$\mathsf{W}(f,i,y)=\sum_{(u,v)\in{E_{i-1,i}}}\left|e_{u}^{T}\mathbf{M}_{i-1\ldots i}y-e_{v}^{T}y\right|$
For every $0\leq l<r\leq n$, the total weight between layer $l$ and layer $r$ of $f$ on $y$ is defined as:
$\mathsf{W}(f,l,r,y)=\sum_{i=l+1}^{r}\mathsf{W}(f,i,\mathbf{M}_{i\ldots r}y).$
We immediately have the following facts:
Lemma E.2.
$0\leq l<m<r\leq n$, $\mathsf{W}(f,l,r,y)=\mathsf{W}(f,l,m,\mathbf{M}_{m\ldots r}y)+\mathsf{W}(f,m,r,y)$
Lemma E.3.
$\mathsf{W}(f)=\mathsf{W}(f,0,n,y_{acc})$, where $y_{acc}$ is the accept vertices vector of $f$.
Proof.
$\mathsf{W}(f,0,n,y_{acc})$ $=\sum_{i=1}^{n}\mathsf{W}(f,i,\mathbf{M}_{i\ldots n}y_{acc})=\sum_{i\in[n]}\sum_{(u,v)\in{E_{i-1,i}}}\left|e_{u}^{T}\mathbf{M}_{i-1\ldots n}y_{acc}-e_{v}^{T}\mathbf{M}_{i\ldots n}y_{acc}\right|$
$=\sum_{(u,v)\in E}\left|\mathbb{E}[f^{u\to}-f^{v\to}]\right|=\mathsf{W}(f)$
We need a main result of [BRRY14].
Lemma E.4 ([BRRY14]).
For every $n,s,w\in\mathbb{N}$ and $\varepsilon>0$, there exists a PRG $G:\{0,1\}^{d}\to(\{0,1\}^{s})^{w}$ such that for every $f\in\mathcal{B}_{n,s,w}$ and every $y\in\mathbb{R}^{w}$:
$\left\|\left(\mathbb{E}_{x\in\{0,1\}^{d}}f^{[l,r]}(G(x))\right)y-\mathbf{M}_{l\ldots r}y\right\|_{\infty}\leq\varepsilon\mathsf{W}(f,l,r,y).$
Furthermore, $d=s+O(\log n(\log\log n+\log(w/\varepsilon))$
We use the error reduction polynomials in [CL24]. Without loss of generality, assume that $n$ is a power of $2$ and define the set $\mathsf{BS}_{n}$ as
$\mathsf{BS}_{n}=\{(l,r)\in[n]^{2}\mid\exists i,k\in\mathbf{N}\cup\{0\},l=i\cdot 2^{k},r=l+2^{k},0\leq l<r\leq n\}.$
Given a set of substochastic matrices $\{\mathbf{M}_{1},\ldots,\mathbf{M}_{n}\}$, suppose for every $(l,r)\in\mathsf{BS}_{n}$, we have a matrix $\mathbf{M}^{(0)}_{l\ldots r}$ that coarsely approximates the product $\mathbf{M}_{l+1}\cdots\mathbf{M}_{r}$. They define the following matrices $\mathbf{M}^{(k)}_{l\ldots r}$ as
$\mathbf{M}^{(k)}_{l\ldots r}=\left\{\begin{array}[]{ll}\mathbf{M}_{r}&\text{if }r=l+1,\\ \sum_{i+j=k}\mathbf{M}^{(i)}_{l\ldots m}\mathbf{M}^{(j)}_{m\ldots r}-\sum_{i+j=k-1}\mathbf{M}^{(i)}_{l\ldots m}\mathbf{M}^{(j)}_{m\ldots r}&\text{otherwise,where }m=(l+r)/2.\end{array}\right.$ (4)
The following result of [CL24] shows that $\mathbf{M}^{(k)}_{l\ldots r}$ is a good approximation as long as $k$ is large enough.
Theorem E.5 (Lemma 3.6 of [CL24]).
For any $0<\varepsilon<\frac{1}{30\log n}$ Assume that for every $(l,r)\in\mathsf{BS}_{n}$
$\forall y\in\mathbb{R}^{w},\left\|\left(\mathbf{M}^{(0)}_{l\ldots r}-\mathbf{M}_{l\ldots r}\right)y\right\|_{\infty}\leq\varepsilon\frac{\mathsf{W}(f,l,r,y)}{\mathsf{W}(f)}$
Then for every $(l,r)\in\mathsf{BS}_{n}$ and $k\in\mathbb{N}$
$\forall y\in\mathbb{R}^{w},\left\|\left(\mathbf{M}^{(k)}_{l\ldots r}-\mathbf{M}_{l\ldots r}\right)y\right\|_{\infty}\leq(30\varepsilon\cdot\log n)^{k}\frac{\mathsf{W}(f,l,r,y)}{\mathsf{W}(f)}$
Now we prove Theorem 3.21
Proof of Theorem 3.21.
Set $k=O(\sqrt{\log(1/\varepsilon)})$. Set $\tau=\frac{\varepsilon^{1/k}}{30\log n\cdot\mathsf{W}(f))}$. Let $\mathsf{PRG}:\{0,1\}^{d}\to(\{0,1\}^{s})^{n}$ be the $\tau$-PRG given by Lemma E.4 and let $(\mathsf{PRG}(\cdot))_{m}$ denote its first $m$ symbols over $\{0,1\}^{s}$. we define $\mathbf{M}^{(0)}_{l\ldots r}:=\mathbb{E}_{x\in\{0,1\}^{d}}f^{[l,r]}((\mathsf{PRG}(x))_{r-l})$ and expand $\mathbf{M}^{(k)}_{0\ldots n}$ as a weighted sum of products of $\mathbf{M}^{(0)}_{l\ldots r}$, we get the following sum:
$\mathbf{M}^{(k)}_{0\ldots n}=\frac{1}{K}\sum_{i\in K}\gamma_{i}\cdot\mathbf{M}^{(0)}_{m_{i,0}\ldots m_{i,1}}\mathbf{M}^{(0)}_{m_{i,1}\ldots m_{i,2}}\ldots\mathbf{M}^{(0)}_{m_{i,t-1}\ldots m_{i,t}},$
where $\gamma_{i}\in\{-K,K\}$ are weights and $0=m_{i,0}\leq m_{i,1}\leq\cdots\leq m_{i,t}=n$ are partitions of $[n]$. We set the $(\log K,K,\varepsilon)$-reduction $(\mathsf{R},\sigma)$ from $\mathcal{R}_{n,s,w}$ to $\mathcal{B}_{t,d,w}$ as
$\mathsf{R}(x_{1},x_{2},\cdots,x_{t},i)$ $=((\mathsf{PRG}(x_{1}))_{m_{i,1}-m_{i,0}},(\mathsf{PRG}(x_{2}))_{m_{i,2}-m_{i,1}},\cdots,(\mathsf{PRG}(x_{t}))_{m_{i,t}-m_{i,t-1}}),$
$\sigma(i)$ $=\gamma_{i}.$
By Lemma E.4, $\forall y\in\mathbb{R}^{w},\left\|\left(\mathbf{M}^{(0)}_{l\ldots r}-\mathbf{M}_{l\ldots r}\right)y\right\|_{\infty}\leq\frac{\varepsilon^{1/k}}{30\log n}\cdot\frac{\mathsf{W}(f,l,r,y)}{\mathsf{W}(f)}$. Then by Theorem E.5, $\left\|\left(\mathbf{M}^{(k)}_{0\ldots n}-\mathbf{M}_{0\ldots n}\right)y_{acc}\right\|_{\infty}\leq\varepsilon\cdot\frac{\mathsf{W}(f,0,n,y_{acc})}{\mathsf{W}(f)}=\varepsilon$, which indicates that our reduction achieves the target precision.
Given the construction of the error reduction polynomial, one can find $K=n^{O(k)}=2^{O(\log n\sqrt{\log(1/\varepsilon)})}$ and $t=O(k\log n)=O(\log n\sqrt{\log(1/\varepsilon)})$. Also, Lemma E.4 indicates $d=s+O(\log n(\log\log n+\log w+\log(\mathsf{W}(f))+\sqrt{\log(1/\varepsilon)})$. That completes the proof.